OpenStudy (anonymous):

Which statement is true for the equation 5n − 4 = 5n − 3? It has infinitely many solutions. It has two solutions. It has one solution. It has no solution.

OpenStudy (lolacole12):

Hi

OpenStudy (gabbyalicorn):

5n−4=5n−3 First you subtract 5n from both sides. 5n−4−5n=5n−3−5n −4=−3 Then you add 4 to both sides. −4+4=−3+4 0=1 Here is your answer: There are no solutions.

OpenStudy (lolacole12):

nun

OpenStudy (lolacole12):

Let's solve your equation step-by-step. 5n−4=5n−3 Step 1: Subtract 5n from both sides. 5n−4−5n=5n−3−5n −4=−3 Step 2: Add 4 to both sides. −4+4=−3+4 0=1 Answer: There are no solutions. same thing

OpenStudy (anonymous):

C

OpenStudy (lolacole12):

yes

OpenStudy (lolacole12):

medal pls

OpenStudy (lolacole12):

no d

OpenStudy (welshfella):

you can also do this just by inspection how can 5x - 4 possibly be equal to 5x - 3 ? there can be no value of x to satisfy that.

OpenStudy (anonymous):

@welshfella so its d?

OpenStudy (lolacole12):

yes d

OpenStudy (lolacole12):

i said that

OpenStudy (welshfella):

definitely